use the function f(x)=3- 4/5x to find the solution to f(x)=0; find the x-intercept of f(x); find the zero of the function f(x)
Answer by stanbon(75887)

use the function f(x)=3- 4/5x to find the solution to f(x)=0;
Solve 3 - (4/5)x = 0
(4/5)x = 3
x = (5/4)3
x = 15/4
x = 3 3/4
------------------------
find the x-intercept of f(x)
x = 3 3/4 is the x-intercept
------------------------
find the zero of the function f(x)
the x-intercept is the zero of f(x)
====================
Cheers,
Stan H.